OpenStudy (anonymous):

For a right circular cylinder, if the radius is tripled and the height is shrunk down to half its size, then what is the surface area if the original radius is 7 inches and the original height is 17 inches?

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

Can you please help me?

OpenStudy (anonymous):

so what is the surface area originally? better yet, what's the formula?

OpenStudy (anonymous):

Im looking

OpenStudy (anonymous):

2 π r2 + 2 π r h

OpenStudy (jdoe0001):

surface area for a right circular cylinder \[2\pi*r*h\] 3times the "r" and half(1/2) the "h" would be \[2\pi*3r*h(\frac{1}{2})\] r=7, h=17 \[2\pi*3(7)*17(\frac{1}{2})\ \]

OpenStudy (anonymous):

1120.98?

OpenStudy (jdoe0001):

well, I get 1121.548

OpenStudy (anonymous):

3892.43 both of you aren't using the total surface area. just the lateral surface area ...

OpenStudy (anonymous):

also known as you forgot to add the circle on the top and bottom and are only using the surfacea area of the side. total surface area = ( 2*Pi*r^2 + 2*Pi*r*h)

OpenStudy (jdoe0001):

right

OpenStudy (jdoe0001):

hmmm, I guess I need a better formulas sheet :|

OpenStudy (mertsj):